WHAT IS THE INTERIOR ANGLE MEASURE DEGREE IN A 11-GON
: WHAT IS THE INTERIOR ANGLE MEASURE DEGREE IN A 11-GON

Answer by stanbon(19015)

WHAT IS THE INTERIOR ANGLE MEASURE DEGREE IN A 11-GON
--------------------
The sum of the interior angles of an 11-gon is (11-2)*180 = 1620 degrees
------------------------
Each of the interior angles is 9*180/11 = 1620/11 = 147.2727... degrees
------------------------
Cheers,
Stan H.